OpenStudy (anonymous):

The perimeter of a rectangle is 40 ft. The length is 6 feet longer than the width. Find the dimensions. Write a system of linear equations and solve the resulting system. Let x be the length and y be the width. Write the first equation 2x+2y= Write the second equation x=y+ What is the length? What is the width?

OpenStudy (anonymous):

some one help me???? Please!!!

OpenStudy (radar):

Let l = length Let w = width we will now write the equation for the perimeter 40=2l +2w Do you follow that for that equation?

OpenStudy (anonymous):

yes

OpenStudy (radar):

Now for the 2nd equation: We were told that l=w + 6

OpenStudy (anonymous):

right

OpenStudy (radar):

Let us put the two equations down one below the other and simplified; 2l + 2w = 40 l - w = 6 Do you want to solve using substitution or the "elimination" method?

OpenStudy (anonymous):

elimination

OpenStudy (anonymous):

so it would be l=w=34?

OpenStudy (anonymous):

l-w=34

OpenStudy (radar):

After reading the instructions in your problem, I should of used x and y but I believe you still understant. The length is 6 ft longer than the width so it is l-w=6

OpenStudy (anonymous):

ok

OpenStudy (radar):

Multiply that equation by 2 2(l-w=6) you will now get 2l - 2w = 12 we now rack em and stack em. 2l + 2w = 40 2l -2w =12 --------------- Lets add 4l =52 l=13 ft Then width will be 6 ft shorter or 7 ft. Check Check 13 + 13 + 7 + 7= 40 Check good 7+6 = 13 check good Bingo good answer.

OpenStudy (radar):

That was the elimination method.

OpenStudy (radar):

If you let l be x and w be y you will have complied with the instructions in your problem.

OpenStudy (anonymous):

ok thank you.